OpenStudy (anonymous):

PreCalculus help anyone? The problem involves me creating an equation so hold on.

OpenStudy (anonymous):

Let \[\log _{b}A=3 \log _{b}C=2 \log _{b}D=5\] What is the value of \[\log _{b}(A^5C^2/D^6)\]?